OpenStudy (anonymous):

6+2c=15 @StudyGurl14 please help

OpenStudy (anonymous):

C=4.5

OpenStudy (kl0723):

isolate your variable: 2c=15-6 2c=9 c=9/2

OpenStudy (anonymous):

that last one is that C=9 divide 2 or.

OpenStudy (kl0723):